OpenStudy (anonymous):

Factor completely: 10x^2 + 2x − 8

OpenStudy (anonymous):

10x2 + 2x − 8= 10x^2+10x-8x-8= 10x*(x+1)-8*(x+1)= (x+1)*(10x-8) Please tell me if you have any doubt about any step.

OpenStudy (anonymous):

that's not one of the answe choices

OpenStudy (anonymous):

answer*

OpenStudy (anonymous):

(x+1)*(10x-8)=2*(x+1)*(5x-4)

OpenStudy (anonymous):

What about that one?

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

You are welcome!

OpenStudy (anonymous):

10x^2+2x-8 2(5x^2+x-4) 2(5x^2+5x-4x-4) 2(5x{x+1}-4{x+1}) 2({5x-4}{x+1})

OpenStudy (anonymous):

thank you too !

OpenStudy (anonymous):

you r welcum =)